# Do Schwartz Higher-Order Values Help Sentence-Level Human Value Detection? When Hard Gating Hurts

### v1

@@ Abstract
Sentence-level human value detection is typically framed as multi-label classification over Schwartz values, but it remains unclear whether Schwartz higher-order (HO) categories provide usable structure. We study this under a strict compute-frugal budget (single 8 GB GPU) on ValueEval’24 / ValuesML (74K English sentences). We compare (i) direct supervised transformers, (ii) HO values pipelines that enforce the hierarchy with hard masks, and (iii) Presence HO values cascades, alongside low-cost add-ons (lexica, short context, topics), label-wise threshold tuning, small instruction-tuned LLM baselines (10B), QLoRA, and simple ensembles. HO categories are learnable from single sentences (e.g., the easiest bipolar pair reaches Macro-), but hard hierarchical gating is not a reliable win: it often reduces end-task Macro- via error compounding and recall suppression. In contrast, label-wise threshold tuning is a high-leverage knob (up to Macro-), and small transformer ensembles provide the most consistent additional gains (up to Macro-). Small LLMs lag behind supervised encoders as stand-alone systems, yet can contribute complementary errors in cross-family ensembles. Overall, HO structure is useful descriptively, but enforcing it with hard gates hurts sentence-level value detection; robust improvements come from calibration and lightweight ensembling.
@@
Morality detection , Human values , Schwartz value theory , Sentence-level classification , Transformer models , Ensemble learning , Large language models[upv]organization=PRHLT Research Center, Universitat Politècnica de València, city=Valencia, postcode=46022, country=Spain
[uev]organization=School of Science, Engineering and Design, Universidad Europea de Valencia, city=Valencia, postcode=46010, country=Spain
[valgrai]organization=Valencian Graduate School and Research Network of Artificial Intelligence (ValgrAI)
@@ 1 Introduction
Human values are enduring guiding principles that shape what people consider important, desirable, or worth protecting (Rokeach, 1973; Schwartz, 1992; Bardi and Schwartz, 2003). Because values are often expressed implicitly in language, detecting them in text matters for computational social science and NLP tasks that analyze public discourse, persuasion, stance, framing, and argumentation at scale (Lazer et al., 2009). Recent surveys of computational morality and value modeling summarize key approaches (lexicon-based signals, supervised classifiers, LLM-centric methods) and emphasize persistent challenges such as contextual ambiguity and domain sensitivity (Reinig et al., 2024). Related work on moral language in political and social media discourse shows that moral/value cues are typically sparse, indirect, and context-dependent (Haidt and Joseph, 2004; Johnson and Goldwasser, 2018).
@@
Among social-science value frameworks, Schwartz’s theory is widely adopted and empirically validated (Schwartz, 1992). The refined theory defines 19 basic values and groups them into higher-order (HO) categories (e.g., Openness to Change vs. Conservation) that capture compatibilities and conflicts (Schwartz, 2012). This hierarchy suggests a potential inductive bias for predicting fine-grained values when labels are sparse or ambiguous. Figure 1 shows the circular motivational continuum of the 19 values, where adjacent values are compatible and opposing values are in conflict (Schwartz et al., 2012).
@@
Recent shared tasks operationalize value detection as sentence-level, multi-label prediction, enabling controlled comparisons (Kiesel et al., 2023). The Touché 2024 Human Value Detection task (ValueEval’24) is built on ValuesML, where sentences are annotated for which Schwartz values they express and whether a value is attained or constrained (The ValuesML Team, 2024; Touché, 2024). Recent work highlights machine learning challenges and benchmark-driven evaluation, showing how dataset design and label distributions shape performance (Rink et al., 2025). These benchmarks reveal realistic difficulties: a sentence may express none, one, or many values; evidence is often implicit and lexically diffuse; and label prevalence is highly imbalanced (Kiesel et al., 2023). These properties strain standard multi-label pipelines and make calibration decisions (thresholds, probability reliability) especially important (Tsoumakas and Katakis, 2007; Guo et al., 2017; Silva Filho et al., 2023).
@@
A natural question is whether Schwartz’s HO structure improves fine-grained sentence-level detection. Hierarchical classification can help by injecting structure, constraining hypotheses, and sharing statistical strength (Silla and Freitas, 2011). But in noisy sentence-level settings, hard constraints can amplify upstream errors: if a parent prediction is uncertain, strict gating can suppress true positives and reduce recall on already sparse labels. This tension is tied to calibration, since hierarchical pipelines are sensitive to thresholds and probability miscalibration (Valmadre, 2022).
@@
We address this tension by evaluating when HO categories help and how to integrate them. Under a compute-frugal setting, we compare: (i) direct multi-label prediction, (ii) a CategoryValues hierarchy that constrains fine-grained outputs, and (iii) a PresenceCategoryValues cascade that first filters sentences predicted to contain any value. We also test low-cost levers that often matter in practice, including threshold calibration and simple ensembling (Wolpert, 1992; Breiman, 1996; Freund and Schapire, 1997; Breiman, 2001). Finally, we benchmark compact instruction-tuned LLMs under the same budget, motivated by evidence that prompting and instruction tuning can be competitive without task-specific architectural changes (Brown et al., 2020; Ouyang et al., 2022; Chung et al., 2024).
Research questions
@@
A focused study of whether Schwartz HO categories improve sentence-level value detection on ValueEval’24/ValuesML, including analyses by canonical bipolar HO pairs.
A comparison of HO-aware strategies (conditioning, hard gating, cascades) that quantifies when hierarchy helps and when it fails due to error propagation.
A compute-frugal evaluation of practical improvements—threshold tuning, calibration-aware decisions, and small ensembles—that yield gains without increasing model size or training cost.
Overall, we provide a practical blueprint for using value structure in NLP pipelines, and show that hard hierarchical constraints are brittle in this sentence-level setting, while calibration and small ensembles deliver the most reliable gains under a compute-frugal budget.
@@ 2.1 Human values and moral frameworks in NLP
Human values are commonly operationalized as relatively stable guiding principles that shape preferences and judgments (Rokeach, 1973; Schwartz, 1992; Bardi and Schwartz, 2003). For an NLP-oriented synthesis of how morality/value constructs are used in text analysis, see Reinig et al. (2024). Among taxonomies, Schwartz’s theory is especially attractive for NLP because it provides (i) a refined, fine-grained set of basic values and (ii) a principled HO organization that captures compatibilities and conflicts (Schwartz et al., 2012). Computational work on moral language often draws on Moral Foundations Theory (MFT) to study moral rhetoric in political and social discourse (Graham et al., 2009). Although MFT and Schwartz address different constructs, both highlight a key modeling challenge: moral/value signals in text are often indirect, sparse, and diffuse rather than explicitly labeled (Graham et al., 2011; Haidt, 2012). Beyond classification, NLP has also been used to elicit structured value representations for downstream systems, such as extracting value promotion schemes (García-Rodríguez et al., 2025).
An extensive line of work explores feature- and lexicon-driven prediction as an interpretable, low-cost alternative to heavy end-to-end models (Hopp et al., 2021; Hoover et al., 2020). Araque et al. (2020) propose MoralStrength, which extends the Moral Foundations Dictionary with embedding-based similarity. González-Santos et al. (2023) study moral foundations assignment in the movie domain using word embeddings and semantic similarity. These studies motivate our compute-frugal perspective: lightweight signals can be useful, but performance depends on how prior structure is injected.
Recent work expands the scope of value detection beyond classic lexicon settings to online community discourse and multimodal platforms, and proposes methods for large text collections. For example, value expressions are analyzed in online communities (Borenstein et al., 2025) and in multimodal influencer content (Starovolsky-Shitrit et al., 2025), while context-dependent markup schemes are proposed for large-scale value/sentiment detection in social media corpora (Rink et al., 2024). In parallel, LLM-based value identification has emerged as a lightweight alternative for text-only settings (Zhu et al., 2025).
@@ 2.2 Benchmarks and shared tasks for value detection
The field has converged on shared benchmarks to enable controlled comparisons and expose realistic difficulty factors such as multi-label outputs, class imbalance, and cross-domain variation. For argumentation, Kiesel et al. (2022) introduce a value-annotated benchmark for identifying human values behind arguments, and the Touché/ValueEval line of tasks systematizes evaluation and reporting (Mirzakhmedova et al., 2024; Kiesel et al., 2023). In Touché 2024, the Human Value Detection task frames value detection as sentence-level prediction under operational constraints typical of applied NLP pipelines (The ValuesML Team, 2024; Touché, 2024).
Within this context, recent systems emphasize cascaded decision processes and threshold control under label sparsity. The best-performing English system reported for Touché/CLEF 2024 uses a cascade to structure decisions and reduce spurious positives (Yeste et al., 2024). Yeste and Rosso (2026) study sentence-level value detection with moral presence gating and compute-frugal transformer ensembles, providing a strong baseline that we extend by focusing on Schwartz HO categories and their use as hierarchical structure.
Beyond shared tasks, recent resource suites aim for broader coverage across frameworks and domains; for example, MoVa aggregates multiple labeled datasets and benchmarks across moral/value theories to enable more generalizable evaluation (Chen et al., 2025).
@@ 2.3 Hierarchical structure and multi-label learning for value prediction
Our core question—whether HO structure helps basic-value prediction—connects to two mature ML literatures: multi-label learning and hierarchical classification. Multi-label classification is difficult when labels are non-mutually exclusive, skewed, and supported by limited positive evidence (Zhang and Zhou, 2014). Hierarchical classification studies how label taxonomies can share statistical strength and impose structure. These insights motivate our HO-aware variants: HO labels may reduce effective sparsity, but rigid enforcement can increase error propagation when parent predictions are uncertain (Silla and Freitas, 2011).
A related theme is the role of context when a single sentence provides limited evidence. Hierarchical document models (e.g., sentence-then-document encoders) improve classification by modeling multi-granular context (Yang et al., 2016). This suggests HO structure may be most effective as guidance combined with careful control of context, rather than as a brittle hard constraint.
@@ 2.4 Calibration, thresholding, and ensemble robustness under imbalance
Because value detection is multi-label and imbalanced, performance is often dominated by decision rules that map scores to binary labels. Classical calibration work (e.g., Platt scaling; Platt (1999)) and later studies on probability reliability show how miscalibration can distort precision/recall trade-offs, especially with per-label thresholds (Silva Filho et al., 2023). This motivates our emphasis on threshold tuning as a compute-frugal but high-leverage component.
Ensembling is another long-standing route to robustness under limited data and noisy supervision (Dietterich, 2000; Rokach, 2010). In value detection, different models can capture complementary cues, so small ensembles often deliver consistent gains without increasing single-model capacity. We build on this principle and the compute-frugal ensemble methodology in Yeste and Rosso (2026) to test whether HO-aware modeling improves beyond calibration and modest diversity.
@@ 2.5 Transformers and instruction-tuned LLMs for moral/value classification
Modern value detection systems typically rely on transformer encoders fine-tuned for classification (Devlin et al., 2019; He et al., 2021). Instruction-tuned LLMs have popularized prompt-based classification as a lightweight alternative that avoids task-specific architectural changes (Zhao et al., 2021; Wei et al., 2022; Liu et al., 2023). Work comparing prompting and supervised adaptation for human values motivates treating prompting vs. fine-tuning as an explicit design choice (Sun, 2024). For sentence-level, sparse multi-label settings, prompt-based LLMs still face calibration and recall challenges, often with less control over score distributions. Parameter-efficient methods such as QLoRA offer a middle ground between pure prompting and full fine-tuning (Hu et al., 2022; Dettmers et al., 2023).
Related work also examines how LLMs represent and align with human values, providing datasets and measurement frameworks. Examples include mapping LLM outputs to Schwartz value dimensions and constructing value-labeled data (Yao et al., 2024), measuring contextual alignment between human and LLM values across scenarios (Shen et al., 2025), and psychometric-style measurement of human/LLM values from text (Ye et al., 2025). Recent analyses probe value consistency and cultural alignment in LLMs (Rozen et al., 2025; Segerer, 2025; Biedma et al., 2024).
Taken together, these threads motivate the design space evaluated in this paper. We adopt the shared-task framing and compute-frugal discipline established in The ValuesML Team (2024); Touché (2024); Yeste and Rosso (2026), and we focus the comparison on how HO structure is injected (conditioning vs. hard gating/cascades) relative to strong, practically motivated baselines based on calibration and small ensembles. This sets up the methodological choices introduced next (Section 3).

### v2

# 1]organization=PRHLT Research Center, Universitat Politècnica de València, city=Valencia, postcode=46022, country=Spain 2]organization=School of Science, Engineering and Design, Universidad Europea de Valencia, city=Valencia, postcode=46010, country=Spain 3]organization=Valencian Graduate School and Research Network of Artificial Intelligence (ValgrAI), city=Valencia, country=Spain

@@ Abstract
Human value detection from single sentences is a sparse, imbalanced multi-label task. We study whether Schwartz higher-order (HO) categories help this setting on ValueEval’24 / ValuesML (74K English sentences) under a compute-frugal budget. Rather than proposing a new architecture, we compare direct supervised transformers, hard HOvalues pipelines, PresenceHOvalues cascades, compact instruction-tuned large language models (LLMs), QLoRA, and low-cost upgrades such as threshold tuning and small ensembles. HO categories are learnable: the easiest bipolar pair, Growth vs. Self-Protection, reaches Macro-. The most reliable gains come from calibration and ensembling: threshold tuning improves Social Focus vs. Personal Focus from to (), transformer soft voting lifts Growth from to , and a Transformer+LLM hybrid reaches on Self-Protection. In contrast, hard hierarchical gating does not consistently improve the end task. Compact LLMs also underperform supervised encoders as stand-alone systems, although they sometimes add useful diversity in hybrid ensembles. Under this benchmark, the HO structure is more useful as an inductive bias than as a rigid routing rule.
@@
[…]
Hard HO gating does not reliably improve end-task value detection.
Threshold tuning yields consistent gains under compute frugal settings.
Small ensembles give the most reliable improvements across HO slices.
Small LLMs lag alone but add diversity in cross family ensembles.
@@ 1 Introduction
Human values are enduring guiding principles that shape what people consider important, desirable, or worth protecting (Rokeach1973; Schwartz1992; BardiSchwartz2003). Because values are often expressed implicitly in language, detecting them in text matters for computational social science and NLP tasks that analyze public discourse, persuasion, stance, framing, and argumentation at scale (Lazer2009). Recent surveys of computational morality and value modeling summarize key approaches (lexicon-based signals, supervised classifiers, LLM-centric methods) and emphasize persistent challenges such as contextual ambiguity and domain sensitivity (Reinig2024). Related work on moral language in political and social media discourse shows that moral/value cues are typically sparse, indirect, and context-dependent (HaidtJoseph2004; JohnsonGoldwasser2018).
@@
Among social-science value frameworks, Schwartz’s theory is widely adopted and empirically validated (Schwartz1992). The refined theory defines 19 basic values and groups them into higher-order (HO) categories (e.g., Openness to Change vs. Conservation) that capture compatibilities and conflicts (Schwartz2012overview). This hierarchy suggests a potential inductive bias for predicting fine-grained values when labels are sparse or ambiguous. Figure 1 shows the circular motivational continuum of the 19 values, where adjacent values are compatible and opposing values are in conflict (Schwartz2012).
@@
Recent shared tasks operationalize value detection as sentence-level, multi-label prediction, enabling controlled comparisons (Kiesel2023). The Touché 2024 Human Value Detection task (ValueEval’24) is built on ValuesML, where sentences are annotated for which Schwartz values they express and whether a value is attained or constrained (ValueEval24Zenodo; Touche2024). Recent work highlights machine learning challenges and benchmark-driven evaluation, showing how dataset design and label distributions shape performance (Rink2025). These benchmarks reveal realistic difficulties: a sentence may express none, one, or many values; evidence is often implicit and lexically diffuse; and label prevalence is highly imbalanced (Kiesel2023). These properties strain standard multi-label pipelines and make calibration decisions (thresholds, probability reliability) especially important (TsoumakasKatakis2007; Guo2017; SilvaFilho2023).
In practical terms, reliable value detection could support several real-world uses. One is value-aware monitoring of political and advocacy messaging, where analysts may want to track how issues are framed in terms of security, tradition, autonomy, or concern for others across campaigns, debates, and media streams (JohnsonGoldwasser2018). A second is large-scale public-opinion and framing analysis, where value signals can help characterize how online communities or large text collections emphasize different normative priorities over time (Borenstein2025; Rink2024). A third is value auditing for AI-mediated communication: as LLMs increasingly generate or transform public-facing text, value detection can serve as a lightweight diagnostic for checking which values their outputs tend to foreground, suppress, or align with (Yao2024; Shen2025; Ye2025).
@@
A natural question is whether Schwartz’s HO structure improves fine-grained sentence-level detection. Hierarchical classification can help by injecting structure, constraining hypotheses, and sharing statistical strength (SillaFreitas2011). But in noisy sentence-level settings, hard constraints can amplify upstream errors: if a parent prediction is uncertain, strict gating can suppress true positives and reduce recall on already sparse labels. This tension is tied to calibration, since hierarchical pipelines are sensitive to thresholds and probability miscalibration (Valmadre2022).
@@
We address this tension through a controlled, compute-bounded empirical study of when HO categories help and how they should be used. Rather than proposing a new hierarchical learner, we compare: (i) direct multi-label prediction, (ii) a CategoryValues hierarchy that constrains fine-grained outputs, and (iii) a PresenceCategoryValues cascade that first filters sentences predicted to contain any value. We also test low-cost levers that often matter in practice, including threshold calibration and simple ensembling (Wolpert1992; Breiman1996; FreundSchapire1997; Breiman2001). Finally, we benchmark compact instruction-tuned LLMs under the same budget, motivated by evidence that prompting and instruction tuning can be competitive without task-specific architectural changes (Brown2020; Ouyang2022; Chung2024). The scope is deliberately benchmark-driven: the goal is to characterize behavior on ValueEval’24 / ValuesML under fixed compute, not to claim immediate generalization across domains, languages, or annotation schemes.
Operationally, this makes the study a set of component-wise ablations under fixed compute, where hierarchy injection, thresholding, auxiliary features, prompting/fine-tuning choices, and ensembling are varied while the evaluation protocol remains constant.
Research questions.
@@
Working hypotheses.
The framework is guided by three testable hypotheses. H1: HO categories should be easier to predict than fine-grained values because they aggregate multiple basic values into coarser labels, reducing sparsity. H2: hard hierarchical gating should improve structural consistency and may improve precision, but can reduce end-task recall through upstream false negatives. H3: under strong label imbalance, label-wise threshold tuning should be a more reliable source of gains than hard gating because it adjusts decision rules without introducing additional upstream failure points.
@@
A careful, compute-bounded empirical study of whether Schwartz HO categories improve sentence-level value detection on ValueEval’24/ValuesML, including analyses by canonical bipolar HO pairs.
A controlled comparison of HO-aware strategies (conditioning, hard gating, cascades) that identifies when hierarchy helps and when it fails due to error propagation.
Evidence that, under this benchmark, calibration-aware thresholding and small ensembles are more reliable sources of gain than hard hierarchical routing.
Overall, we contribute a practical empirical characterization of when value structure helps under fixed compute, and show that hard hierarchical constraints are brittle in this sentence-level setting, while calibration and small ensembles deliver the most reliable gains under this benchmark.
@@ 2.1 Human values and moral frameworks in NLP
Human values are commonly operationalized as relatively stable guiding principles that shape preferences and judgments (Rokeach1973; Schwartz1992; BardiSchwartz2003). For an NLP-oriented synthesis of how morality/value constructs are used in text analysis, see Reinig2024. Among taxonomies, Schwartz’s theory is especially attractive for NLP because it provides (i) a refined, fine-grained set of basic values and (ii) a principled HO organization that captures compatibilities and conflicts (Schwartz2012). Computational work on moral language often draws on Moral Foundations Theory (MFT) to study moral rhetoric in political and social discourse (Graham2009). Although MFT and Schwartz address different constructs, both highlight a key modeling challenge: moral/value signals in text are often indirect, sparse, and diffuse rather than explicitly labeled (Graham2011; Haidt2012). Beyond classification, NLP has also been used to elicit structured value representations for downstream systems, such as extracting value promotion schemes (GarciaRodriguez2025).
An extensive line of work explores feature- and lexicon-driven prediction as an interpretable, low-cost alternative to heavy end-to-end models (Hopp2021; Hoover2020). Araque2020 propose MoralStrength, which extends the Moral Foundations Dictionary with embedding-based similarity. GonzalezSantos2023 study moral foundations assignment in the movie domain using word embeddings and semantic similarity. These studies motivate our compute-frugal perspective: lightweight signals can be useful, but performance depends on how prior structure is injected.
Recent work expands the scope of value detection beyond classic lexicon settings to online community discourse and multimodal platforms, and proposes methods for large text collections. For example, value expressions are analyzed in online communities (Borenstein2025) and in multimodal influencer content (Starovolsky-Shitrit2025), while context-dependent markup schemes are proposed for large-scale value/sentiment detection in social media corpora (Rink2024). In parallel, LLM-based value identification has emerged as a lightweight alternative for text-only settings (Zhu2025).
@@ 2.2 Benchmarks and shared tasks for value detection
The field has converged on shared benchmarks to enable controlled comparisons and expose realistic difficulty factors such as multi-label outputs, class imbalance, and cross-domain variation. For argumentation, Kiesel2022 introduce a value-annotated benchmark for identifying human values behind arguments, and the Touché/ValueEval line of tasks systematizes evaluation and reporting (Mirzakhmedova2024; Kiesel2023). In Touché 2024, the Human Value Detection task frames value detection as sentence-level prediction under operational constraints typical of applied NLP pipelines (ValueEval24Zenodo; Touche2024).
Within this context, recent systems emphasize cascaded decision processes and threshold control under label sparsity. The best-performing English system reported for Touché/CLEF 2024 uses a cascade to structure decisions and reduce spurious positives (Yeste2024). Yeste2026 study sentence-level value detection with moral presence gating and compute-frugal transformer ensembles, providing a strong baseline that we extend by focusing on Schwartz HO categories and their use as hierarchical structure.
Beyond shared tasks, recent resource suites aim for broader coverage across frameworks and domains; for example, MoVa aggregates multiple labeled datasets and benchmarks across moral/value theories to enable more generalizable evaluation (Chen2025).
@@ 2.3 Hierarchical structure and multi-label learning for value prediction
Our core question—whether HO structure helps basic-value prediction—connects to two mature ML literatures: multi-label learning and hierarchical classification. Multi-label classification is difficult when labels are non-mutually exclusive, skewed, and supported by limited positive evidence (ZhangZhou2014). Hierarchical classification studies how label taxonomies can share statistical strength and impose structure. These insights motivate our HO-aware variants: HO labels may reduce effective sparsity, but rigid enforcement can increase error propagation when parent predictions are uncertain (SillaFreitas2011; Wang2024).
A related theme is the role of context when a single sentence provides limited evidence. Hierarchical document models (e.g., sentence-then-document encoders) improve classification by modeling multi-granular context (Yang2016), and prior analyses in hierarchical text classification show that representation choices across levels can materially affect downstream performance (Stein2019). This suggests HO structure may be most effective as guidance combined with careful control of context, rather than as a brittle hard constraint.
@@ 2.4 Calibration, thresholding, and ensemble robustness under imbalance
Because value detection is multi-label and imbalanced, performance is often dominated by decision rules that map scores to binary labels. Classical calibration work (e.g., Platt scaling; Platt1999) and later studies on probability reliability show how miscalibration can distort precision/recall trade-offs, especially with per-label thresholds (SilvaFilho2023). This motivates our emphasis on threshold tuning as a compute-frugal but high-leverage component.
Ensembling is another long-standing route to robustness under limited data and noisy supervision (Dietterich2000; Rokach2010). In value detection, different models can capture complementary cues, so small ensembles often deliver consistent gains without increasing single-model capacity. This aligns with recent dynamic ensemble work for multi-label classification under label dependence and imbalance (Zhu2023). We build on this principle and the compute-frugal ensemble methodology in Yeste2026 to test whether HO-aware modeling improves beyond calibration and modest diversity. More specifically, whereas prior work mainly motivates calibration and ensembling as generally useful tools under imbalance, our contribution is to contrast them directly with hard hierarchical routing under a fixed benchmark and compute budget, showing that the former are the more reliable source of gains in this setting.
@@ 2.5 Transformers and instruction-tuned LLMs for moral/value classification
Modern value detection systems typically rely on transformer encoders fine-tuned for classification (Devlin2019; He2021). Instruction-tuned LLMs have popularized prompt-based classification as a lightweight alternative that avoids task-specific architectural changes (Zhao2021; Wei2022; Liu2023). Work comparing prompting and supervised adaptation for human values motivates treating prompting vs. fine-tuning as an explicit design choice (Sun2024). For sentence-level, sparse multi-label settings, prompt-based LLMs still face calibration and recall challenges, often with less control over score distributions. Parameter-efficient methods such as QLoRA offer a middle ground between pure prompting and full fine-tuning (Hu2022; Dettmers2024).
Recent LLM work usefully splits into two strands. The first is closest to our task: value detection/classification from text, where LLMs are used to assign value labels to inputs. Recent examples include EAVIT, which studies LLM-based human value identification from text, and MoVa, which targets more generalizable morals/value classification across a broader aggregated benchmark suite (Zhu2025; Chen2025). The second studies value alignment/representation in the models themselves, asking which values LLMs express, prioritize, or align with rather than predicting values in input text. Examples include Value FULCRA, which maps LLM outputs to Schwartz value dimensions (Yao2024), UniVaR, which learns a high-dimensional representation of value distributions in LLMs across models and languages (Cahyawijaya2025), ValueCompass, which measures contextual alignment between human and LLM values across scenarios (Shen2025), and psychometric-style measurement of human/LLM values from text (Ye2025). Recent analyses further probe value consistency and cultural alignment in LLMs (Rozen2025; Segerer2025; Biedma2024). Our experiments are positioned in the first strand, while the second mainly motivates why value-aware analysis of LLM outputs matters.
Taken together, these threads motivate the design space evaluated in this paper. We adopt the shared-task framing and compute-frugal discipline established in ValueEval24Zenodo; Touche2024; Yeste2026, and we focus the comparison on how HO structure is injected (conditioning vs. hard gating/cascades) relative to strong, practically motivated baselines based on calibration and small ensembles. This sets up the methodological choices introduced next (Section 3).